### v2

# Probabilistic Pontryagin’s Maximum Principle for Continuous-Time Model-Based Reinforcement Learning

Abstract
Without exact knowledge of the true system dynamics, optimal control of non-linear continuous-time systems requires careful treatment of epistemic uncertainty. In this work, we propose a probabilistic extension to Pontryagin’s maximum principle by minimizing the mean Hamiltonian with respect to epistemic uncertainty. We show minimization of the mean Hamiltonian is a necessary optimality condition when optimizing the mean cost, and propose a multiple shooting numerical method scalable to large-scale probabilistic dynamical models, including ensemble neural ordinary differential equations. Comparisons against state-of-the-art methods in online and offline model-based reinforcement learning tasks show that our probabilistic Hamiltonian formulation leads to reduced trial costs in offline settings and achieves competitive performance in online scenarios. By bridging optimal control and reinforcement learning, our approach offers a principled and practical framework for controlling uncertain systems with learned dynamics.
I Introduction
Optimal decision-making under uncertainty is a fundamental challenge in learning-based control, particularly when the system dynamics are unknown and have to be learned from limited and noisy data. Achieving data-efficiency in such settings is crucial, as real-world experiments can incur significant costs. Moreover, effectively handling epistemic uncertainty, which arises from insufficient data, is critical, as it directly influences robustness of control, as well as the exploitation/exploitation balance during learning.
A common approach for decision making under uncertainty involves minimizing the mean cost with respect to the posterior distribution of system dynamics. However, optimizing this mean cost can become challenging in practice, especially when the posterior distribution is multi-modal or has significant variance. Consequently, black-box stochastic optimization methods are typically employed to mitigate local minima, but these ignore the inherent structure of the optimal control problem [1].
Pontryagin’s maximum principle (PMP) provides a different perspective. It characterizes optimal solutions indirectly by describing the necessary conditions of optimality, which implicitly defines the optimal control inputs [2]. As one of the cornerstones in optimal control theory, PMP has driven significant advances, particularly in deterministic control settings [3]. By leveraging the inherent optimality conditions of the control problem, PMP enables computationally efficient approaches to optimization and provides insights into the underlying structure of optimal solutions.
Related work. Principles from optimal control theory serve as foundational tools in learning-based control, providing guidelines for optimal action selection under uncertainty. Several optimal control methodologies have been extended into uncertain or learning-driven contexts, notably the iterative linear quadratic regulator (iLQR) [4] and path integral (PI) control [5]. PMP, originally formulated for deterministic systems, has also seen interest in robust control [6] and, more recently, learning-based settings [7]. An open-loop variant of this was further explored in [8]. However, these prior PMP-based methods are formulated for deterministic system dynamics and do not address epistemic uncertainty arising from limited or noisy data.
Our work closely aligns with Bayesian control paradigms and Bayesian reinforcement learning (RL), where the design of optimal policies explicitly incorporates uncertainty [9]. A seminal work is the probabilistic inference for learning and control (PILCO) framework [10], which learns discrete-time Gaussian process models of system dynamics and efficiently computes policy gradients through closed-form Gaussian approximations. Extensions of Bayesian methods to model predictive control (MPC) have demonstrated efficacy by planning under uncertain system dynamics, leveraging PMP conditions for propagating Gaussian uncertainty over a moving horizon [11]. Model-based planning under uncertainty has also significantly advanced model-based reinforcement learning (MBRL), where methods like probabilistic ensembles with trajectory sampling (PETS) [12] efficiently tackle planning under uncertainty in high-dimensional tasks.
Our work also intersects with recent advances in continuous-time system identification and learning-based control [13], where discrete-time measurements are used to reconstruct the continuous-time dynamics that characterize many physical and biological systems. Extensions into RL have investigated continuous-time policies [14] and efficient exploration [15].
Contribution. Our contributions are threefold. Firstly, we propose a probabilistic extension of classical PMP by minimizing a mean Hamiltonian to explicitly incorporate epistemic uncertainty, and we demonstrate that this provides a necessary optimality condition for control under uncertain dynamics. This probabilistic formulation enhances robustness and performance in scenarios with incomplete or noisy model knowledge. Secondly, we introduce a practical numerical implementation for the mean Hamiltonian approach using a forward multiple shooting method combined with moving horizon optimization, specifically tailored to probabilistic ordinary differential equation (ODE) models. Lastly, we experimentally validate our approach, showing improved or competitive performance compared to state-of-the-art methods in both offline and online RL scenarios, highlighting the value of our method for efficient learning and robust decision-making.
II Problem setting
We study a continuous-time deterministic differential equation $\boldsymbol{f}$, whose solution is given by:
$\boldsymbol{x}(t)=\boldsymbol{x}_{0}+\int_{t_{0}}^{t_{f}}\boldsymbol{f}\left(\boldsymbol{x}(\tau),\boldsymbol{u}(\tau),\tau\right)\operatorname{d}\!{\tau}$ (II.1)
where $\boldsymbol{x}_{0}\in\mathcal{X}\subset\mathds{R}^{d_{x}}$ is the initial state and $\boldsymbol{u}\colon[t_{0},t_{f}]\to\mathcal{U}\subset\mathds{R}^{d_{u}}$ is the control input. The objective is to find the optimal control input $\boldsymbol{u}^{*}$, that is, the input that minimizes the cost function:
$C(\boldsymbol{u})$ $=\int_{t_{0}}^{t_{f}}L\bigl{(}\boldsymbol{x}(\tau),\boldsymbol{u}(\tau)\bigr{)}\operatorname{d}\!{\tau}+\Phi\left(\boldsymbol{x}(t_{f})\right)$ (II.2)
where $\Phi(\boldsymbol{x})$ is the termination cost associated with ending in state $\boldsymbol{x}(t_{f})$ at time $t_{f}$ and $L(\boldsymbol{x},\boldsymbol{u})$ is the state cost associated with taking action $\boldsymbol{u}$ in state $\boldsymbol{x}$. Both the termination cost and state cost are assumed differentiable with respect to their inputs. We are interested in solving the following optimal control problem:
$\boldsymbol{u}^{*}(t)$ :-$\coloneq\operatorname*{argmin}_{\boldsymbol{u}\in\mathcal{U}}C(\boldsymbol{u}):$ (II.3)
[…]
$\boldsymbol{x}(t_{0})$ $=\boldsymbol{x}_{0}\enspace.$
Although the true function $\boldsymbol{f}$ is unknown, it can be evaluated by interacting with the system and collecting data over the course of $N$ trials with time horizon $t_{f}$. Let $\mathcal{D}_{n}\coloneq\{(t^{n}_{i},\boldsymbol{y}^{n}_{i},\boldsymbol{u}^{n}_{i})\}$ denote the data collected in trial $n$. Here, $t^{n}_{i}\in[t_{0},t_{f}]$ is the time point corresponding to observation $i$ in trial $n$, $\boldsymbol{x}_{i}^{n}\coloneq\boldsymbol{x}(t^{n}_{i})$ is the true system state at time $t^{n}_{i}$ in trial $n$, $\boldsymbol{y}^{n}_{i}\sim\mathcal{N}(\boldsymbol{x}_{i}^{n},\sigma_{n}^{2}\boldsymbol{I})$ is a noisy measurement of the state with i.i.d. Gaussian noise, and $\boldsymbol{u}^{n}_{i}$ is the applied control input corresponding to observation $i$ for trial $n$. It is assumed that the true control function can be reconstructed for $t\in[t_{0},t_{f}]$ using interpolation techniques.
In MBRL, a dynamical model is learned based on previous interactions with the system. In the online variant, this dataset consists initially of only a single trial generated by a random controller. Then iteratively, a model is trained on the available data, the optimal action is determined using the model, and the newly observed trajectory is added to the data.
II-A Probabilistic dynamical models
When learning a dynamical model from limited data, it is crucial to employ efficient exploration strategies that separate epistemic uncertainty, associated with reducible uncertainty due to lack of knowledge of the true system, from aleatoric uncertainty, associated with irreducible uncertainty due to inherent randomness such as sensor or process noise. A general approach for quantifying uncertainty is to assign a prior distribution over the parameters and a likelihood to the possibly noisy observations, and use Bayes’ theorem to marginalize over the uncertainty in model parameters:
$\boldsymbol{\theta}$ $\sim p(\boldsymbol{\theta})$ (II.4)
$p(\mathcal{D}\mid\boldsymbol{\theta})$ $=\prod_{i=1}^{T}\mathcal{N}\left(\boldsymbol{y}_{i}\mid\boldsymbol{x}(t_{i};\boldsymbol{\theta}),\sigma^{2}\boldsymbol{I}\right)$ (II.5)
$p(\boldsymbol{\theta}\mid\mathcal{D})$ $\propto p(\mathcal{D}\mid\boldsymbol{\theta})\>p(\boldsymbol{\theta})$ (II.6)
where $T$ is the number of discrete-time observations throughout $[t_{0},t_{f}]$ and $\boldsymbol{\theta}\in\Theta=\mathbb{R}^{p}$ are the parameters. For a given realization $\boldsymbol{\theta}$, the following continuous-time model is defined:
$\dot{\boldsymbol{x}}(t;\boldsymbol{\theta})$ $=\boldsymbol{f}\bigl{(}\boldsymbol{x}(t;\boldsymbol{\theta}),\boldsymbol{u}(t),t;\boldsymbol{\theta})$ (II.7)
where $\boldsymbol{f}$ is differentiable in its first two arguments. The posterior distribution can then be estimated using Bayesian sampling methods, such as Hamiltonian Monte Carlo (HMC).
II-B Mean cost objective
A common objective in MBRL is the expected cost over the parameter distribution of the dynamical model, which reflects the epistemic uncertainty:
$\bar{C}(\boldsymbol{u})\coloneq\mathbb{E}_{\boldsymbol{\theta}}\left[C(\boldsymbol{u};\boldsymbol{\theta})\right]$ (II.8)
where
$C(\boldsymbol{u};\boldsymbol{\theta})$ $=\int_{0}^{t_{f}}L\bigl{(}\boldsymbol{x}(\tau;\boldsymbol{\theta}),\boldsymbol{u}(\tau)\bigr{)}\operatorname{d}\!{\tau}+\Phi\bigl{(}\boldsymbol{x}(t_{f};\boldsymbol{\theta})\bigr{)}$
and $\mathbb{E}_{\boldsymbol{\theta}}$ is shorthand for $\mathbb{E}_{\boldsymbol{\theta}\sim p(\boldsymbol{\theta}\mid\mathcal{D})}$. The cost functional $C(\boldsymbol{u};\boldsymbol{\theta})$ denotes the cost for realization $\boldsymbol{\theta}$. This results in the optimal control problem:
$\boldsymbol{u}^{*}(t)$ :-$\coloneq\operatorname*{argmin}_{\boldsymbol{u}\in\mathcal{U}}\bar{C}(\boldsymbol{u}):$ (II.9)
[…]
$\boldsymbol{x}(t_{0};\boldsymbol{\theta})$ $=\boldsymbol{x}_{0}\enspace.$
III Probabilistic control via PMP
III-A Pontryagin’s maximum principle
The mean cost objective can be optimized via the direct method or via the indirect method. In the former, the control function is discretized and optimized as a real-valued vector, whereas in the latter, first the necessary optimality conditions for the continuous-time optimal control problem are described using PMP and then the control function is discretized. A key function is the Hamiltonian:
$\mathcal{H}\bigl{(}\boldsymbol{x},\boldsymbol{\lambda},\boldsymbol{u},t\bigr{)}\coloneq L\bigl{(}\boldsymbol{x},\boldsymbol{u}\bigr{)}+\boldsymbol{\lambda}^{\top}\boldsymbol{f}\bigl{(}\boldsymbol{x},\boldsymbol{u},t\bigr{)}$
where $\mathcal{H}\colon\mathcal{X}\times\mathbb{R}^{d_{x}}\times\mathcal{U}\times\mathbb{R}\to\mathbb{R}$ and $\boldsymbol{\lambda}\in\mathbb{R}^{d_{x}}$ are the co-states (or adjoints). For a given function $\boldsymbol{u}(t)$, the state and costate equations are given as:
$\dot{\boldsymbol{x}}(t)$ $=\boldsymbol{f}\bigl{(}\boldsymbol{x}(t),\boldsymbol{u}(t),t\bigr{)}$ (III.1)
$\dot{\boldsymbol{\lambda}}(t)$ $=-\nabla_{\boldsymbol{x}}\mathcal{H}\bigl{(}\boldsymbol{x}(t),\boldsymbol{\lambda}(t),\boldsymbol{u}(t),t\bigr{)}$ (III.2)
$\boldsymbol{\lambda}(t_{f})$ $=\nabla\Phi\bigl{(}\boldsymbol{x}(t_{f})\bigr{)}$ (III.3)
for $t\in[t_{0},t_{f}]$, where the notation $(\nabla_{y}f)_{i,j}\coloneq\partial f_{j}/{\partial y_{i}}$ denotes Jacobians, in this case $\nabla_{\boldsymbol{x}}\mathcal{H}=\left(\frac{\partial\mathcal{H}}{\partial x_{1}},\dots,\frac{\partial\mathcal{H}}{\partial x_{d_{x}}}\right)^{\top}$ and $\nabla\Phi=\left(\frac{\partial\Phi}{\partial x_{1}},\dots,\frac{\partial\Phi}{\partial x_{d_{x}}}\right)^{\top}$. Equation (III.3) is commonly known as the transversality condition, and provides a boundary condition for the co-states. If $\boldsymbol{u}^{*}(t)$ is an optimal solution to $C(\boldsymbol{u})$, the following necessary condition of optimality is derived via PMP:
$\boldsymbol{u}^{*}(t)=\operatorname*{argmin}_{\boldsymbol{u}}\mathcal{H}\left(\boldsymbol{x}^{*}(t),\boldsymbol{\lambda}^{*}(t),\boldsymbol{u},t\right)$ (III.4)
where $\boldsymbol{x}^{*}(t)$ and $\boldsymbol{\lambda}^{*}(t)$ satisfy conditions (III.1) – (III.3). We will refer to equation (III.4) as the minimum principle. The indirect approach provides candidate solutions for the optimal control $\boldsymbol{u}^{*}(t)$ that satisfy the necessary optimality conditions for the cost functional. Next, it is detailed how these conditions extend to the probabilistic setup.
III-B Probabilistic Hamiltonian
The posterior distribution over system dynamics $p(\boldsymbol{\theta}\mid\mathcal{D})$ in Eq. (II.6) implicitly defines a distribution over Hamiltonian functions $p(\mathcal{H}\mid\mathcal{D})$ through its dependence on the dynamics. Sampling a realization from the dynamics posterior yields a corresponding Hamiltonian:
$\mathcal{H}(\boldsymbol{x},\boldsymbol{\lambda},\boldsymbol{u},t;\boldsymbol{\theta})\coloneq L\left(\boldsymbol{x},\boldsymbol{u}\right)+\boldsymbol{\lambda}^{\top}\boldsymbol{f}\left(\boldsymbol{x},\boldsymbol{u},t;\boldsymbol{\theta}\right).$ (III.5)
Our goal is to find a control function $\boldsymbol{u}^{*}(t)$ that minimizes the mean cost functional $\bar{C}(\boldsymbol{u})$. For a single realization $\boldsymbol{\theta}$, the necessary optimality conditions for $C(\boldsymbol{u};\boldsymbol{\theta})$ follow from the minimum principle:
$\boldsymbol{u}^{*}(t;\boldsymbol{\theta})=\operatorname*{argmin}_{\boldsymbol{u}}\mathcal{H}\left(\boldsymbol{x}^{*}(t;\boldsymbol{\theta}),\boldsymbol{\lambda}^{*}(t;\boldsymbol{\theta}),\boldsymbol{u},t;\boldsymbol{\theta}\right)$
with corresponding state and costate trajectories $\boldsymbol{x}^{*}(t;\boldsymbol{\theta})$ and $\boldsymbol{\lambda}^{*}(t;\boldsymbol{\theta})$ satisfying Eqs. (III.1) – (III.3). For the necessary optimality conditions of $\bar{C}(\boldsymbol{u})$, we consider the mean minimum principle:
$\boldsymbol{u}^{*}(t)=\operatorname*{argmin}_{\boldsymbol{u}}\mathbb{E}_{\theta}\left[\mathcal{H}\left(\boldsymbol{x}(t;\boldsymbol{\theta}),\boldsymbol{\lambda}(t;\boldsymbol{\theta}),\boldsymbol{u},t;\boldsymbol{\theta}\right)\right]$ (III.6)
where $\boldsymbol{x}(t;\boldsymbol{\theta})$ and $\boldsymbol{\lambda}(t;\boldsymbol{\theta})$ are the state and costate trajectories that satisfy Eqs. (III.1) – (III.3) for a realization $\boldsymbol{\theta}\sim p(\boldsymbol{\theta}\mid\mathcal{D})$. These optimality conditions result in a family of state and co-state trajectories, that are optimal in expectation but not for each individual realization of $\boldsymbol{\theta}$. Despite this, the co-states $\boldsymbol{\lambda}(t;\boldsymbol{\theta})$ remain interpretable as cost sensitivity by satisfying the transversality condition.
III-C Necessary optimality conditions for the mean cost
We now consider if the mean minimum principle is a necessary condition for optimality of the mean cost $\bar{C}(\boldsymbol{u})$. We take a variational approach that assumes unbounded controls, but note that the constrained setting can be approached using the needle (or spike) variation of the proof [16].
Theorem 1 (Mean Pontryagin Minimum Principle).
Let $\boldsymbol{u}^{*}(t)$ be an optimal control that minimizes the mean cost functional $\bar{C}(\boldsymbol{u})$, and let $\mathcal{U}=\mathbb{R}^{d_{u}}$. Then a necessary condition for optimality is:
[…]
Proof.
We assume that $\mathcal{H}$ is differentiable following $\boldsymbol{f}$ and $L$’s differentiability, and that $\nabla_{\boldsymbol{u}}\mathcal{H}$ is integrable w.r.t. $p(\boldsymbol{\theta}\mid\mathcal{D})$. We consider the Gateaux derivative of the mean cost:
$\delta\bar{C}(\boldsymbol{u})=\lim\limits_{\epsilon\to 0}\frac{\bar{C}(\boldsymbol{u}+\epsilon\delta\boldsymbol{u})-\bar{C}(\boldsymbol{u})}{\epsilon}$
where $\delta\boldsymbol{u}(t)$ is an arbitrary variation defined over $[t_{0},t_{f}]$. The first variation can be obtained by isolating the first order terms around the optimal solution $\boldsymbol{u}^{*}(t)$:
$\bar{C}(\boldsymbol{u}^{*}+\epsilon\delta\boldsymbol{u})-\bar{C}(\boldsymbol{u}^{*})\approx\delta\bar{C}\big{|}_{\boldsymbol{u}^{*}(t)}(\delta\boldsymbol{u})\epsilon\enspace.$
Since by definition $\boldsymbol{u}^{*}(t)$ is optimal for $\bar{C}$, its first variation must be zero.
We first consider the deterministic case for a realization $\boldsymbol{\theta}$, where $\boldsymbol{u}^{*}(t;\boldsymbol{\theta})$ is optimal for $C(\boldsymbol{u};\boldsymbol{\theta})$. A well-known result in optimal control theory states that the first variation is given as [3, Eq. (5.1.13)]:
$\delta C_{\boldsymbol{\theta}}\big{|}_{\boldsymbol{u}^{*}(t;\boldsymbol{\theta})}(\delta\boldsymbol{u})=\int^{t_{f}}_{t_{0}}\bigl{[}$ $\nabla_{\boldsymbol{u}}\mathcal{H}_{*}\bigr{]}\delta\boldsymbol{u}(t)\operatorname{d}\!{t}\enspace,$
where $\mathcal{H}_{*}=\mathcal{H}\bigl{(}\boldsymbol{x}^{*}(t;\boldsymbol{\theta}),\boldsymbol{\lambda}^{*}(t;\boldsymbol{\theta}),\boldsymbol{u}^{*}(t;\boldsymbol{\theta}),t;\boldsymbol{\theta}\bigr{)}$ and $C_{\boldsymbol{\theta}}=C(\boldsymbol{u};\boldsymbol{\theta})$. By the fundamental lemma of calculus of variations, it follows that:
$\nabla_{\boldsymbol{u}}\mathcal{H}_{*}=0,\quad\text{for }t\in[t_{0},t_{f}]\enspace.$
Using this result, we now turn to the mean cost. Under standard regularity conditions (via dominated convergence), we interchange the derivative and expectation:
$\delta\bar{C}(\boldsymbol{u})=\delta\mathbb{E}_{\boldsymbol{\theta}}[C(\boldsymbol{u};\boldsymbol{\theta})]=\mathbb{E}_{\boldsymbol{\theta}}[\delta C(\boldsymbol{u};\boldsymbol{\theta})]\enspace.$
Using the earlier result of the first variation:
$\mathbb{E}_{\boldsymbol{\theta}}[\delta C_{\boldsymbol{\theta}}\big{|}_{\boldsymbol{u}^{*}(t)}(\delta\boldsymbol{u})]$
$=\mathbb{E}_{\boldsymbol{\theta}}\bigl{[}\int^{t_{f}}_{t_{0}}\nabla_{\boldsymbol{u}}\mathcal{H}\bigl{(}\boldsymbol{x}(t;\boldsymbol{\theta}),\boldsymbol{\lambda}(t;\boldsymbol{\theta}),\boldsymbol{u}^{*}(t),t;\boldsymbol{\theta}\bigr{)}\delta\boldsymbol{u}\operatorname{d}\!{t}\bigr{]}$
$=\int^{t_{f}}_{t_{0}}\nabla_{\boldsymbol{u}}\mathbb{E}_{\boldsymbol{\theta}}\bigl{[}\mathcal{H}\bigl{(}\boldsymbol{x}(t;\boldsymbol{\theta}),\boldsymbol{\lambda}(t;\boldsymbol{\theta}),\boldsymbol{u}^{*}(t),t;\boldsymbol{\theta}\bigr{)}\bigr{]}\delta\boldsymbol{u}\operatorname{d}\!{t}$
[…]
using the assumption that $\nabla_{\boldsymbol{u}}\mathcal{H}$ is integrable w.r.t. $p(\boldsymbol{\theta}\mid\mathcal{D})$. By the fundamental lemma of the calculus of variations, it follows that:
[…]
This shows that the optimal control $\boldsymbol{u}^{*}(t)$ is only required to satisfy the optimality conditions in the expectation of $p(\boldsymbol{\theta}\mid\mathcal{D})$, extending the minimum principle to probabilistic dynamical models that capture epistemic uncertainty.
IV Numerical methods for probabilistic Hamiltonian
IV-A Trajectory sampling
Computing the mean over Hamiltonians $\bar{\mathcal{H}}$ in Eq. (III.6) requires estimating the uncertainty in state and costate trajectories $\boldsymbol{x}^{*}(t;\boldsymbol{\theta})$ and $\boldsymbol{\lambda}^{*}(t;\boldsymbol{\theta})$ over $t\in[t_{0},t_{f}]$. We adopt the trajectory sampling technique to capture this variability introduced by the dynamics posterior [12]. The mean Hamiltonian is approximated by aggregating the sampled Hamiltonian functions:
$\boldsymbol{u}^{*}(t)\approx\operatorname*{argmin}_{\boldsymbol{u}}\frac{1}{M}\sum_{i=1}^{M}\mathcal{H}\left(\boldsymbol{x}(t;\boldsymbol{\theta}_{i}),\boldsymbol{\lambda}(t;\boldsymbol{\theta}_{i}),\boldsymbol{u},t;\boldsymbol{\theta}_{i}\right)$ (IV.1)
where $\boldsymbol{\theta}_{i}\sim p(\boldsymbol{\theta}\mid\mathcal{D})$ for $i=1,\dots,M$, and each conditioned Hamiltonian is computed along the trajectories $\boldsymbol{x}(t;\boldsymbol{\theta}_{i})$ and $\boldsymbol{\lambda}(t;\boldsymbol{\theta}_{i})$, which comply with the PMP conditions. An alternative formulation is to consider the mean of the optimal controls $\bar{\boldsymbol{u}}^{*}(t)=\mathbb{E}_{\boldsymbol{\theta}}[\boldsymbol{u}^{*}(t;\boldsymbol{\theta})]$:
$\bar{\boldsymbol{u}}^{*}(t)\approx\frac{1}{M}\sum^{M}_{i=1}\boldsymbol{u}^{*}(t;\boldsymbol{\theta}_{i})\enspace.$ (IV.2)
By not evaluating the mean Hamiltonian with a shared optimal control input for the mean cost, the mean optimal control approximation is only a valid necessary optimality condition for the mean cost when the Hamiltonian function is convex. By uncoupling the optimization process between different realizations of the dynamics posterior, however, the solutions can be computed in parallel, therefore leading to improved computational complexity. Both approaches leverage Monte Carlo estimates of the state and co-state trajectories to efficiently evaluate the $M$ samples using GPU parallelization.
IV-B Two-point boundary value problem
For each realization $\boldsymbol{\theta}_{i}$, the initial state $\boldsymbol{x}(t_{0};\boldsymbol{\theta}_{i})=\boldsymbol{x}_{0}$ and the final co-state $\boldsymbol{\lambda}(t_{f};\boldsymbol{\theta}_{i})=\nabla_{\boldsymbol{x}}\Phi\bigl{(}\boldsymbol{x}(t_{f};\boldsymbol{\theta}_{i})\bigr{)}$ are shared. This gives rise to a two-point boundary value problem (TPBVP) that satisfies the coupled differential equations $\dot{\boldsymbol{x}}(t;\boldsymbol{\theta}_{i})=\boldsymbol{f}\bigl{(}\boldsymbol{x}(t;\boldsymbol{\theta}_{i}),\boldsymbol{\lambda}(t;\boldsymbol{\theta}_{i})\bigr{)}$ and $\dot{\boldsymbol{\lambda}}(t;\boldsymbol{\theta}_{i})=-\nabla_{\boldsymbol{x}}\mathcal{H}(\boldsymbol{x}(t;\boldsymbol{\theta}_{i}),\boldsymbol{\lambda}(t;\boldsymbol{\theta}_{i}),\boldsymbol{u}(t),t;\boldsymbol{\theta}_{i})$, where the initial value of the co-state $\boldsymbol{\lambda}(t_{0};\boldsymbol{\theta}_{i})$ is unknown. Due to the continuous-time formulation, numerical ODE solvers with adaptive step-size solvers can be leveraged to improve stability and control the numerical error.
We solve this using the forward method [17], which guesses an initial co-state $\boldsymbol{\lambda}(t_{0};\boldsymbol{\theta}_{i})$, integrates the states and co-states forward, and adjusts the guess based on the termination errors $\boldsymbol{\lambda}(t_{f};\boldsymbol{\theta}_{i})-\nabla_{\boldsymbol{x}}\Phi\bigl{(}\boldsymbol{x}(t_{f};\boldsymbol{\theta}_{i})\bigr{)}$. Root-finding methods (e.g., Newton’s method) or least-squares optimization (e.g., Levenberg-Marquardt) refine the guess iteratively. For robustness with non-linear dynamics for long time horizons, we employ the multiple shooting formulation [17]. This method breaks down the long time horizon into multiple shorter intervals and enforces continuity conditions between these, to improve convergence and stability.
The mean Hamiltonian approach shares controls across dynamics function realizations, leading to $MSd_{x}$ decision variables, where $S$ is the number of shooting segments. In contrast, the mean optimal control variant decouples the optimization process of the sample realizations by only averaging after the control function is optimized. This results in $M$ independent optimization problems with $Sd_{x}$ decision variables, which can be solved in parallel.
IV-C Model predictive control
The TPBVP describes an open-loop trajectory optimization problem. To obtain a closed-loop control, principles from non-linear MPC can be leveraged by solving the trajectory optimization at every time point under a moving horizon of $t\in[t_{0},t_{0}+H]$ [17]. The optimized control function is then applied until a new system measurement becomes available, and the optimization process is repeated.
V Experimental Results
We demonstrate our approach on three classical control problems. Code implementing the experiments is available at https://github.com/DavidLeeftink/probabilistic-pontryagin-control.
V-A Van der Pol stabilization
We first evaluate our method on stabilizing a Van der Pol oscillator, a non-linear dynamical system commonly used in electrical circuit design and computational neuroscience [18]. We consider the following control formulation:
$\dot{x}_{1}$ $=\mu\left(x_{2}-\frac{1}{3}x_{1}^{3}-x_{1}\right)$
$\dot{x}_{2}$ $=-x_{1}+u$
where $\mu=1.5$. The goal is stabilization of the system at the unstable fixed point at the origin, under a quadratic cost function. This requires precise control actions to prevent repulsion towards the attracting limit cycle.
We formulate an offline RL task by generating a a dataset of $n=25$ trajectories (length $T=10$) with the Van der Pol system, whose dynamics are assumed unknown, under random periodic control inputs (Schroeder sweeps [19]) from uniformly sampled initial conditions. We then use this data to model the dynamics probabilistically, by training an ensemble of neural ODEs which capture the epistemic uncertainty. The resulting probabilistic dynamical model is then used to construct a controller by either following the probabilistic Pontryagin approaches or by directly optimizing the mean cost.
To explicitly investigate the effect of epistemic uncertainty modeling, we compare our approach using both deterministic and probabilistic dynamical models trained on identical datasets. We benchmark the mean Hamiltonian (PMP-Mean $\mathcal{H}$) and mean posterior (PMP-Mean $u^{*}$) formulations against direct optimization methods: Sequential Quadratic Programming (SQP), Broyden–Fletcher–Goldfarb–Shanno (BFGS) [20], improved Cross-Entropy Method (iCEM) [21, 22], and Adam [23]. We provide further experimental details in the Appendix, Table III.
TABLE I: Comparison of optimization methods for Van der Pol stabilization with true and learned dynamics (mean cost $\pm$ std. dev. over 15 datasets).
Method True Model NODE Prob. NODE
PMP-Mean $\mathcal{H}$ 9.85 12.71 $\pm$ 2.45 11.57 $\pm$ 1.73
PMP-Mean $u^{*}$ 10.59 $\pm$ 0.88
iCEM 10.08 13.90 $\pm$ 2.80 14.06 $\pm$ 2.81
SQP 10.04 13.10 $\pm$ 2.52 13.69 $\pm$ 2.14
BFGS 10.03 13.13 $\pm$ 2.59 13.67 $\pm$ 2.75
Adam 10.05 14.03 $\pm$ 2.66 14.27 $\pm$ 3.03
Table I summarizes the performance in terms of accumulated cost over 15 repetitions. The indirect PMP approach achieves improved cost performance over its direct counterpart, under both the deterministic and probabilistic dynamical model. The lowest cost is achieved under the probabilistic Pontryagin approaches, where the mean optimal control approximation slightly outperforms the mean Hamiltonian approach, suggesting that explicitly modeling the epistemic uncertainty improves performance.
V-B Cart pole swing up
TABLE II: Comparison of optimization methods for the Cart pole control problem with true and learned dynamics (mean cost $\pm$ std. dev. over 15 datasets).
Method True Model NODE Prob. NODE
PMP-Mean $\mathcal{H}$ 12.57 112.22 $\pm$ 126.99 36.78 $\pm$ 31.45
PMP-Mean $u^{*}$ 64.44 $\pm$ 77.95
iCEM 7.23   87.55 $\pm$ 109.56 109.51 $\pm$ 183.34
SQP 12.57 111.52 $\pm$ 110.37   77.98 $\pm$ 107.10
BFGS 12.57 147.95 $\pm$ 168.82   81.78 $\pm$ 132.58
Adam 12.89 251.31 $\pm$ 270.43 311.77 $\pm$ 349.75
Next, we consider a cart pole swing-up task, where the objective is to balance the pendulum upright from an initial down position under quadratic cost [24]:
$\ddot{x}$ $=\frac{lm\sin(\theta)\dot{\theta}^{2}+u+m\,g\cos(\theta)\sin(\theta)}{M+m(1-\cos^{2}(\theta))}$
[…]
We design an offline RL task in similar fashion to the Van der Pol experiment, where the dynamics of the cart pole are assumed to be unknown and have to be learned from observed trajectories. We generate $n=50$ trajectories using random controls and uniformly sampled initial conditions, which serve as data. This results in sparse observations near the goal state. We compare the performance of the previously mentioned trajectory optimization methods on deterministic and probabilistic neural ODEs. Further experimental details are described in the Appendix, Table III.
Figure 1: Comparison of optimization methods for the cart pole control task with deterministic and probabilistic dynamical models. The boxplot shows the distribution of cost over 15 datasets, including the median, interquartile range, and outliers. Due to its high cost average, the Adam optimizer is not visualized.
Table II and Figure 1 present the trial costs averaged over 15 datasets, comparing deterministic and probabilistic dynamical models across various optimization methods. Under limited available data, adopting the probabilistic dynamical model improves performance for most optimizers, emphasizing the role of capturing epistemic uncertainty in learning-based control tasks. The mean Hamiltonian formulation achieves the lowest mean and median trial costs out of the considered optimization methods, outperforming direct methods. In contrast, the mean posterior approximation, theoretically suited only for convex Hamiltonians, performs worse with higher variability and average cost.
V-C Duffing oscillator stabilization
Lastly, we consider an online RL study involving the Duffing oscillator, a non-linear system modeling a mass in a double well potential, which can be regarded as a model of a steel beam that is deflected towards two magnets [25]:
$\dot{x}_{1}$ $=x_{2}$
$\dot{x}_{2}$ $=-\delta x_{2}-\alpha x_{1}-\beta x_{1}^{3}+\gamma u$
Under periodic forcing, chaotic behavior can be observed. In our configuration, the system has two stable attractors, separated by an unstable fixed point at the origin. The objective is to drive the system to the unstable fixed point.
In the online RL setting, one starts from a single random-policy trial, and iteratively trains the probabilistic dynamical model, updates the controller, and augments the dataset after each trial. We compare the mean Hamiltonian and mean posterior methods to the state-of-the-art PETS algorithm [12], which combines probabilistic dynamical models with MPC and the (i)CEM optimizer. We adapt PETS to the continuous-time setting using probabilistic neural ODEs as dynamical models. Further experimental details are described in the Appendix, Table III.
Figure 2: Comparison of mean Hamiltonian PMP approach with PETS, for online RL of a Duffing oscillator (cost $\pm$ std. dev. over 10 independent repetitions).
Figure 2 presents the average trial costs over 10 sequential iterations (averaged over 10 independent repetitions), with the dashed reference line indicating performance under known dynamics. Both PETS and the mean Hamiltonian approach converge to similar final costs, although PETS shows slight cost oscillations. The mean posterior approximation initially performs better but later becomes unstable, likely due to the inadequate aggregation of independently optimized control functions in the presence of multiple attractors near the goal state.
VI Conclusion and discussion
We have proposed a probabilistic extension of PMP that incorporates epistemic uncertainty by minimizing the mean of Hamiltonian functions with respect to a posterior distribution over system dynamics. This approach enables the application of classical PMP techniques to probabilistic dynamical models, which commonly arise in learning-based control and RL settings, particularly under limited data.
Our numerical implementation combines forward multiple shooting with moving horizon optimization. This method indirectly optimizes the state and co-state trajectories, implicitly defining continuous-time optimal control functions. This naturally aligns with the continuous-time nature of many real-world control systems, while still only requiring discrete-time observations. Although we adopted neural ODE ensembles as a probabilistic dynamical model, the proposed framework generalizes easily to other probabilistic dynamical models, including Bayesian neural ODEs or systems with known functional forms but uncertain parameters.
Our experiments demonstrate the practical effectiveness of our approach in offline RL tasks under limited data availability, including the Van der Pol oscillator and a cart pole system. Comparisons against direct trajectory optimization methods, such as SQP and (i)CEM, show that our proposed mean Hamiltonian formulation achieves improved trial costs. Our results further underline the advantage of explicitly modeling the epistemic uncertainty: Under identical neural architectures, the probabilistic ensemble model consistently outperforms its deterministic counterpart, indicating that capturing epistemic uncertainty can lead to improved control performance. Additionally, in online RL settings, such as the stabilization of a Duffing oscillator, the proposed approach achieves competitive trial cost and data efficiency, comparable to the state-of-the-art MBRL method PETS.
Although these results are promising, computational challenges arise in higher-dimensional systems or large ensemble sizes, due to solving the co-state equations via a root-finding task with cubic complexity. However, the sparse Jacobian structure and parallelizable computations of the sampled dynamics provide avenues for future algorithmic improvements or GPU acceleration. Alternatively, the forward-backward iterative approach can be adopted, which avoids the cubic complexity but may require additional iterations until convergence [26].
Future theoretical work could explore extending the optimality conditions to bounded control sets, a common setting for many practical control problems including the experiments considered in this work. From a practical perspective, our probabilistic PMP approach is especially suited for applications requiring precise control under uncertainty on slower timescales, such as personalized treatment planning in healthcare or process control, where state constraints like dosage or safety limits often arise. The PMP framework can be extended by incorporating the state constraints in the Hamiltonian via additional co-state variables.
ACKNOWLEDGMENTS
We would like to thank Dr. Yuzhen Qin for his valuable feedback and insightful comments during the review of this manuscript.
APPENDIX
TABLE III: System parameters and control cost matrices used in the experiments.
System Parameters $Q$ $Q_{f}$ $R$ $\sigma$ $\mathcal{U}$ $x_{0}$ $x^{*}$ $t_{0}$ $t_{f}$ $\Delta t$ (MPC) $H$ Solver
Van der Pol $\mu=1.5$ $\boldsymbol{I}$ $\boldsymbol{I}$ $0.5$ $0.01$ $[-2,2]$ $(1,1)^{\top}$ $(0,0)^{\top}$ 00 $10$ $0.05$ $3$ dopri8
Cart pole $l$ $=m=M=1$ $g$ $=9.81$ $(1,1,0.1,0.1)^{\top}$ $(1,5,1,1)^{\top}$ $50.0$ $[-20,20]$ $(0,0,0,0)^{\top}$ $(1,\pi,0,0)^{\top}$ 00 $5$ $0.02$ $1$ dopri8
Duffing $\alpha$ $=-1,\quad\beta=2$ $\delta$ $=0.2,\quad\gamma=1$ $5\cdot\boldsymbol{I}$ $5\cdot\boldsymbol{I}$ $1$ $0.01$ $[-2,2]$ $(1.5,1)^{\top}$ $(0,0)^{\top}$ 00 $5$ $0.05$ $2$ dopri8
VI-A Experimental details
For all experiments, cubic interpolation is used to construct a continuous-time control function from a discretized control vector, and a 5th-order Dormand-Prince solver is used to solve the states (and co-states for PMP) [27]. The following settings were used for the respective experiments:
Van der Pol
Following the approach described in [12], we consider an ensemble size of $M=5$, with each neural ODE consisting of two hidden layers with 32 units, resulting in parameters $\boldsymbol{\theta}\in\mathbb{R}^{1248}$. Our indirect methods use $S=4$ shooting segments. All trajectory optimization methods employ moving horizon optimization (horizon $H=3$) and 15 iterations per timestep.
Cart pole
The dynamical model architecture is consistent with the previous Van der Pol setup, using ensemble neural ODEs ($\boldsymbol{\theta}\in\mathbb{R}^{1376}$) with identical control interpolation and ODE solver. Here, we use a shorter MPC horizon of $H=1$ with 25 iterations per time step, and two shooting segments ($S=2$) to ensure computational efficiency.
Duffing
We adopt an ensemble of neural ODEs with $M=5$ as a probabilistic dynamical model, consistent with the previous van der Pol setup ($\boldsymbol{\theta}\in\mathbb{R}^{1248}$) with identical control interpolation and ODE solver. We employ MPC with horizon $H=2$ and adjusted integration steps suitable for the Duffing oscillator’s faster time scale, and use $S=5$ shooting segments for the indirect methods.
The system parameters and control cost matrices used in all experiments are described in Table III.